Predetermined Overhead Rate, Applied Overhead, Unit Cost

Ripley, Inc., costs products using a normal costing system. The following data are available for last year:

Budgeted:
      Overhead      $294,400
      Machine hours      87,000
      Direct labor hours      9,200
Actual:
      Overhead      $293,800
      Machine hours      84,800
      Direct labor hours       8,920
      Prime cost      $694,560
      Number of units      140,000

Overhead is applied on the basis of direct labor hours.

Required:

1. What was the predetermined overhead rate?
$___ per direct labor hour

2. What was the applied overhead for last year?
$____

3. Was overhead over- or underapplied, and by how much?
Underapplied overhead  = $___

4. What was the total cost per unit produced? (Carry your answer to four significant digits.)
$____ per unit
Answer #1

1.

Predetermined overhead rate = Estimated overheads/Direct labor hours

= 294,400/9,200

= $32 per direct labor hour

2.

Applied overhead = Actual direct labor hours x Predetermined overhead rate

= 8,920 x 32

= $285,440

3.

Actual overhead = $293,800

Applied overhead = $285,440

Since Applied overhead is less than Actual overhead, hence overhead is under applied.

Under applied overhead  = Actual overhead - Applied overhead

= $293,800 - $285,440

= $8,360

4.

Total cost =   Prime cost + Applied overhead

= 694,560 + 285,440

= $980,000

Total cost per unit = Total cost/Number of units produced

= 980,000/140,000

= $7

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.